40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 13:56:56 2011, kiwamu, Update, ASC, C1ASS : status update 
    Reply  Thu Sep 22 17:48:20 2011, kiwamu, Update, SUS, bad oplev on ETMY OLETMY.png
       Reply  Thu Sep 22 20:12:54 2011, kiwamu, Update, SUS, ETMY oplev whitening engaged OLETMY_WF.png
    Reply  Mon Sep 26 12:41:27 2011, kiwamu, Update, ASC, C1ASS : status update 
       Reply  Fri Sep 30 01:18:39 2011, kiwamu, Update, ASC, C1ASS : status update 
Message ID: 5523     Entry time: Thu Sep 22 20:12:54 2011     In reply to: 5521
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ETMY oplev whitening engaged 

The whitening filters for the ETMY oplevs are back.

The whitening board had been in the rack but the ADC was connected directly to the oplev interface board without going through the whitening board.

In fact the interface board and the whitening board had been already connected. So the ADC was making a shortcut.

I disconnected the ADC from the interface board and plugged it to the output of the whitening board.

Here is an example of the new open-loop transfer function with the whitening filters.

OLETMY_WF.png

 Note :

before the measurement I increased the control gain by an arbitrary number to obtain gain of more than 1 around 1 Hz.

Quote from #5521

I will check the whitening filters.

ELOG V3.1.3-